In a completely randomized design, ten subjects were assigned to each of three treatments of a factor. The partially completed ANOVA table is shown below. Complete parts a through d.

Source DF Sum of Squares Mean Square F-Ratio P-Value

Treatment (Between) 831.84

Error (Within)

Total 1103.41

?a) What are the degrees of freedom for? treatment, error, and? total?
The degrees of freedom for treatment are
nothing.
?(Simplify your? answer.)
The degrees of freedom for error are
nothing.
?(Simplify your? answer.)
The degrees of freedom for the total are
nothing.
?(Simplify your? answer.)
?b) What is? SSE?
SSEequals=nothing
?(Type an integer or a? decimal.)
?c) What is? MSTr?
MSTrequals=277.28
?(Round to two decimal places as? needed.)
?d) What is? MSE?
MSEequals=nothing
?(Round to two decimal places as? needed.)
Answer #1

No. of treatments , k = 3

Number of subjects for each treatment are n = 10

The Total number of subjects = N = nk = 10*3 = 30

ANOVA TABLE

Source of Variation df SS MSS F p value
Treatment (Between) k-1= 3-1 = 2 831.84 831.84/2 =415.92 415.92/10.06 =41.35 0.000
Error (Within) N-k = 30-3=27 1103.41-831.84=271.57 271.57/27 =10.06
Total N-1= 30-1 = 29 1103.41